Suppose x is a normally distributed random variable with μ=46 and σ=3. Find a value x0 of the random variable x that satisfies the following equations or statements.

a.​ 10% of the values of x are less than x0.

b.​ 80% of the values of x are less than x0.

c.​ 1% of the values of x are greater than x0.


1
Expert's answer
2021-07-19T05:43:41-0400

a. "P(Z<z_0)=0.1."

"z_0=-1.28."

"\frac{x_0-46}{3}=-1.28."

"x_0=46-1.28*3=42.16."


b. "P(Z<z_0)=0.8."

"z_0=0.84."

"\frac{x_0-46}{3}=0.84."

"x_0=46+0.84*3=48.52."


c. "P(Z>z_0)=0.01\to P(Z<z_0)=0.99."

"z_0=2.33."

"\frac{x_0-46}{3}=2.33."

"x_0=46+2.33*3=52.99."